OpenStudy (anonymous):

The 20th term of the series 2 *4 + 4*6 + 6*8+...... is?

OpenStudy (anonymous):

\[\huge a_{n}=4n(n+1)\]

OpenStudy (klimenkov):

The general member of sequence is \[2n\cdot(2n+2)\] Let n=20 and get the result

OpenStudy (anonymous):

Wat is that

OpenStudy (anonymous):

is this in AGP

OpenStudy (anonymous):

@satellite73

OpenStudy (klimenkov):

what means AGP?

OpenStudy (anonymous):

Arithmetico - geometric progression

OpenStudy (anonymous):

Plzz help me!

OpenStudy (klimenkov):

Not sure but i think its not

OpenStudy (anonymous):

22*24

OpenStudy (anonymous):

The answer should be 1680

OpenStudy (anonymous):

@satellite73 plzzzz hrlp

OpenStudy (anonymous):

@mukushla @klimenkov ur method is correct but hw

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

oops sorry it should be 40*42 here it is. .. 2n(2n+2)

OpenStudy (anonymous):

hw did u figure out the equation

OpenStudy (callisto):

2 *4 + 4*6 + 6*8+ ... First term = 2 *4 = 2(1) * 2(1+1) Second term = 4*6 = 2(2) * 2(2+1) Third term = 6*8 = 2(3) * 2(3+1) => General term = 2n* 2(n+1)

OpenStudy (anonymous):

thxxx

OpenStudy (callisto):

Welcome :)